### v1

# Can LLM-based Financial Investing Strategies Outperform the Market in Long Run?

Abstract
Large Language Models (LLMs) have recently been leveraged for asset pricing tasks and stock trading applications, enabling AI agents to generate investment decisions from unstructured financial data. However, most evaluations of LLM timing-based investing strategies are conducted on narrow timeframes and limited stock universes, overstating effectiveness due to survivorship and data-snooping biases. We critically assess their generalizability and robustness by proposing FINSABER [^1] Data and code available at https://github.com/waylonli/FINSABER/., a backtesting framework evaluating timing-based strategies across longer periods and a larger universe of symbols. Systematic backtests over two decades and 100+ symbols reveal that previously reported LLM advantages deteriorate significantly under broader cross-section and over a longer-term evaluation. Our market regime analysis further demonstrates that LLM strategies are overly conservative in bull markets, underperforming passive benchmarks, and overly aggressive in bear markets, incurring heavy losses. These findings highlight the need to develop LLM strategies that are able to prioritise trend detection and regime-aware risk controls over mere scaling of framework complexity.
Can LLM-based Financial Investing Strategies Outperform the Market in Long Run?
Weixian Waylon Li1        Hyeonjun Kim2        Mihai Cucuringu3,4        Tiejun Ma1 1School of Informatics, University of Edinburgh 2Global Finance Research Center, Sungkyunkwan University 3Department of Mathematics, University of California Los Angeles 4Department of Statistics, University of Oxford {waylon.li, tiejun.ma}@ed.ac.uk
1 Introduction
Large language models (LLMs) are increasingly used in financial decision-making, especially for generating investment actions such as Buy, Hold, or Sell (Ding et al., 2024a; Fatouros et al., 2025). These so-called LLM timing-based investing strategies leverage LLMs’ ability to interpret historical and real-time data to autonomously trade. From sentiment-driven trading Zhang et al. (2024a) to sophisticated multi-agent systems Yu et al. (2023); Zhang et al. (2024b), a growing body of work has explored the potential of LLMs as autonomous financial agents.
Backtesting is the standard method for assessing investment strategies, simulating them on historical data to evaluate profitability and robustness (Chan, 2021). However, current LLM investing research suffers from fragmented, underdeveloped evaluation practices. Most studies assess performance over short periods, on few symbols, and often omit code release, limiting reproducibility. As summarised in Table 1, several recent methods evaluate over under a year, with fewer than ten stocks, and benchmark only against naïve baselines like Buy-and-Hold. Such short horizons and narrow stock universes lead to three well-documented sources of bias: survivorship bias (Garcia and Gould, 1993), where delisted or failed stocks are omitted; look-ahead bias (Chan, 2021), where future information inadvertently influences past decisions; and data-snooping bias (Bailey et al., 2015), where strategy performance is inflated through repeated testing on the same data. These biases can result in misleading performance assessments and undermine the validity of claimed improvements over traditional methods. This raises a central question: Can LLM-based investing strategies survive longer and broader robustness evaluations?
Method Type (Period, Symbols) Code
MarketSenseAI Sentiment Driven (1y3m, 100)
TradingGPT Multi Agents (N/A, N/A)
FinMem Multi Agents (6m, 5)
FinAgent Multi Agents (6m, 6)
FinRobot Multi Agents (N/A, N/A)
TradExpert Multi Agents (1y, 30)
FinCon Multi Agents (8m, 8)
TradingAgents Multi Agents (3m, 3)
MarketSenseAI 2.0 Multi Agents (2y, 100)
Table 1: Summary of current LLM-based investing strategies. Evaluation period and symbol coverage are combined for compactness.
While recent efforts such as Wang et al. (2025) and Hu et al. (2025) have begun to address backtesting of deep learning (DL)-based strategies and benchmarking of LLM-based time-series forecasting, a standard framework for rigorously evaluating LLM investing strategies remains absent. Moreover, these works do not effectively tackle the mitigation of biases in quantitative trading. To fill this gap, we introduce FINSABER, a comprehensive framework for benchmarking LLM timing-based investing strategies that supports longer backtesting periods, a broader and more diverse symbol universe, and explicit bias mitigation. Specifically, our main contributions are:
  1. 1.
    We propose FINSABER, the first comprehensive evaluation framework for LLM-based investing strategies that integrates 20 years of multi-source data, expands symbol coverage via unbiased selection, and mitigates survivorship, look-ahead, and data-snooping biases.
  2. 2.
    We empirically reassess prior claims and show that LLM advantages reported in recent studies often vanish under broader and longer evaluations, indicating that many conclusions are driven by selective or fragile setups.
  3. 3.
    We conduct regime-specific analysis and reveal that LLM strategies underperform in bull markets due to excessive conservatism and suffer disproportionate losses in bear markets due to inadequate risk control.
  4. 4.
    We offer guidance for future LLM strategy design, arguing that regime-awareness and adaptive risk management are more critical than increasing architectural complexity.
Altogether, our work provides empirical guidance for LLM-based investment research, advocating for the development of strategies that are able to adjust to dynamically-changing market conditions.
2 Related Works
Recent work using LLMs as investors directly employ LLMs to make investing decisions (Ding et al., 2024a). The most common approach leverages LLMs’ sentiment analysis capabilities, using either general-purpose LLMs (e.g., GPT, LLaMA (Touvron et al., 2023), Qwen (Bai et al., 2023)) or fine-tuned financial variants like FinGPT (Yang et al., 2023) to generate sentiment scores for trading decisions (Lopez-Lira and Tang, 2023; Wu, 2024; Kirtac and Germano, 2024; Zhang et al., 2024a). However, these approaches stop short of forming complete trading strategies, which require not only directional forecasts, but also realistic liquidity sizing for mitigating impact, development of execution rules for trade timing and risk management, and incorporation of trading costs.
More advanced approaches move beyond sentiment scores by summarising and reasoning over multi-source financial text. For example, Fatouros et al. (2024) introduce a memory module that stores summarised financial data, retrieved during trading to guide decisions. Similarly, LLMFactor (Wang et al., 2024) learns to extract profitable factors from historical news aligned with price movements and applies them to future market forecasts.
A growing body of work incorporates LLM-based agents (Guo et al., 2024), where either one specialised agent or multiple collaborative agents are employed to perform financial analysis or predictions. Notable examples include FinMem (Yu et al., 2023), FinAgent (Zhang et al., 2024b), FinRobot (Yang et al., 2024), TradExpert (Ding et al., 2024b), FinCon (Yu et al., 2024), TradingAgents (Xiao et al., 2024) and MarketSenseAI 2.0 (Fatouros et al., 2025). Some models also incorporate reinforcement learning (RL) for iterative self-improvement (Ding et al., 2023; Koa et al., 2024).
3 Definitions of Investing Strategies
3.1 Timing-based Investing Strategies
A timing-based investing strategy consists of predefined rules or algorithms that dictate buy, sell, or hold decisions for a given asset based on market data. The goal is to systematically anticipate short-term price movements to generate positive profit and loss (PnL) returns.
Let $P_{t}$ denote the adjusted closing price of a stock on day $t$, for $t=1,2,\dots,T$. A timing-based strategy determines a daily position $S_{t}\in\{-1,0,+1\}$, where $+1$ indicates a Buy, $-1$ a Sell, and 00 a Hold. Such strategies are typically driven by signals extracted from technical indicators, macroeconomic variables, sentiment, or machine learning models, aiming to exploit temporal inefficiencies in asset pricing.
3.2 Selection-based Investing Strategies
A selection-based strategy focuses on identifying a subset of assets from a broader universe to include in a portfolio. Unlike timing-based strategies, which focus on entry and exit points, selection-based approaches prioritise choosing the right assets based on their relative expected performance and the magnitude of the forecasts.
Rather than assigning precise portfolio weights, our setup considers equal-weighted or top-$k$ selection, where the aim is to allocate capital across high-conviction assets based on signals such as momentum, volatility, or model predictions. These strategies are periodically refreshed to adapt to changing market conditions.
The goal is to consistently favour assets expected to outperform, enabling long-term alpha through disciplined selection and rebalancing.
4 Why Broader and Longer Evaluation Matters
Robust evaluation of financial strategies demands carefully designed backtests. Unlike typical machine learning tasks with large, clean datasets, financial data is noisy, nonstationary, and limited in scope. As a result, backtests are especially prone to three major sources of bias: survivorship bias, look-ahead bias, and data-snooping bias, each of which can inflate perceived performance and lead to misleading conclusions (Chan, 2021).
Survivorship Bias.
This occurs when backtests include only currently active stocks while ignoring delisted or bankrupt assets. Such omissions systematically overstate returns and understate risk (Joubert et al., 2024). A common cause is using today’s S&P 500 constituents as the historical investment universe. This practice introduces what Garcia and Gould (1993) call “preinclusion bias”, also a form of look-ahead bias where future index membership influences past decisions. The impact is well-documented: Grinblatt and Titman (1989) and Elton et al. (1996) estimate annual return distortions between 0.1% and 0.9%, and Brown et al. (1992) show that even small distortions can misrepresent performance persistence.
Look-ahead Bias.
Look-ahead bias arises when a strategy uses information that would not have been known at the time of decision-making (Chan, 2021). This includes selecting features, parameters, or symbols based on full-period outcomes, thereby introducing future knowledge into the backtest.
Data-snooping Bias.
Also known as multiple testing bias, this occurs when repeated experimentation on the same dataset leads to overfitting. In finance, where sample sizes are small and the signal-to-noise ratio is very low, this bias is particularly problematic. Bailey et al. (2015) showed that evaluating many strategies on overlapping data inflates false positive rates, and that standard hold-out validation techniques often fail to guard against this issue.
The Case for Broader and Longer Evaluation.
Addressing these biases requires evaluating strategies across longer periods and broader asset universes. For daily trading, at least three years of data is generally recommended, while weekly and monthly strategies benefit from 10 to 20 years or more (Bailey et al., 2015). Gatev et al. (2006) tested pairs trading on 40 years of daily data, but Do and Faff (2010) extended this to 48 years and found profitability declined, highlighting the need for long-term evaluation. Likewise, recent deep learning models in finance rely on multi-year datasets to ensure robustness (Feng et al., 2019; Wang et al., 2022; Xia et al., 2024).
Symbol selection is another critical factor. Many LLM-based investing studies selectively use only a small number of well-known stocks such as TSLA and AMZN. These are both historical winners, which limits generalisability and embeds both survivorship and look-ahead bias into the evaluation. Omitting delisted or underperforming stocks distorts performance metrics and presents an incomplete picture of real-world investing conditions.
Therefore, backtests must address survivorship bias, look-ahead bias, and data-snooping bias explicitly. Broader and longer evaluations, using historically accurate stock universes and spanning multiple market regimes, are essential for producing reliable, generalisable results that reflect real investing conditions.
5 FINSABER
As discussed in §4, existing evaluations of LLM-based investors often suffer from survivorship bias, look-ahead bias, and data-snooping bias. These issues are largely due to limited evaluation periods and narrow stock selections. To address these limitations, we propose FINSABER—a backtesting framework for Financial INvesting Strategy Assessment with Bias mitigation, Expanded time, and Range of symbols.
FINSABER is composed of three core modules: a multi-source data module, a strategies base, and a bias-aware backtesting pipeline. Together, they support robust and reproducible benchmarking across four experimental setups described in §6. An overview is shown in Figure 1.
Figure 1: Overview of the FINSABER Backtest Framework. The central pipeline illustrates the backtesting process. The framework includes a Strategies Base Module (green), which covers both selection-based and timing-based strategies, and a Multi-source Data Module (yellow), integrating diverse financial data inputs.
Multi-source Data for LLM Benchmarking.
FINSABER supports benchmarking across a wide range of investing strategies, including those driven by large language models. To do so, it integrates structured and unstructured financial data sources commonly used by LLM-based agents: historical stock prices, financial news, and company filings (10-K, 10-Q), all spanning from 2000 to 2024. To prevent look-ahead bias, all data inputs are aligned with each backtest window using only information available prior to the start date. Delisted stocks are included, and open-source equivalents are provided for reproducibility (more detail in Appendix A).
Modular Strategy Design and Bias-Mitigated Backtest Pipeline.
Strategies in FINSABER are categorised into selection-based (which pick stocks to trade) and timing-based (which decide when to trade) where timing-based strategies base includes rule-based, ML/DL, RL, and LLM investors and can be extended with custom modules (Appendix B). To mitigate survivorship bias, FINSABER uses historical S&P 500 constituent lists available at the beginning of each evaluation window and explicitly includes delisted stocks in the investable universe. To reduce data-snooping bias, the backtest pipeline employs rolling-window evaluation across multiple periods and asset selections. Both the window size and step are customisable, allowing researchers to simulate realistic evaluation settings across diverse market regimes. Timing-based strategies are paired with selection strategies to ensure broader symbol coverage and avoid overfitting to a narrow stock set or short horizon.
Evaluation Metrics.
FINSABER adopts three categories of evaluation metrics: return, risk, and risk-adjusted performance. Return metrics measure profitability and include Annualised Return (AR) and Cumulative Return (CR). Risk metrics quantify uncertainty and downside exposure, including Annualised Volatility (AV) and Maximum Drawdown (MDD). Risk-adjusted metrics assess capital efficiency and include the Sharpe Ratio (SPR) and Sortino Ratio (STR).
High returns alone do not imply strategy quality. Risk-adjusted metrics such as SPR and STR are more informative, especially in finance where capital efficiency and downside risk are critical (Chan, 2021). These metrics are standard in the literature (Cont, 2001; DeMiguel et al., 2007) and widely used in recent LLM-based investing benchmarks (Zhang et al., 2024b; Yu et al., 2024). Formal definitions and formulas are provided in Appendix C.
6 Experiments
Our experiments directly address the methodological flaws in current LLM-based investing evaluations identified in §4, particularly survivorship and data-snooping biases caused by selective stock choices and short evaluation periods. We show how these practices yield overly optimistic results and how FINSABER enables fairer assessment.
Specifically, our experiments are structured into two main parts: (1) Demonstrating pitfalls of selective evaluation: We replicate the originally reported performance results of recent LLM investors on selectively chosen periods and symbols. We then extend this evaluation period while keeping the original symbols unchanged, revealing how performance deteriorates over longer time horizons. (2) Establishing fair and robust comparisons: We leverage a composite approach, incorporating systematic stock-selection methods to explicitly mitigate survivorship and data-snooping biases, thus enabling fair and reliable assessments of LLM investor performance. We only consider go-long positions in trading, as existing LLM strategies do not model or support short-selling behaviour.
6.1 Pitfalls of Selective Evaluation
Revisiting Originally Reported Claims.
We begin by replicating prior setups from the literature, particularly evaluating the FinMem and FinAgent strategy on TSLA, NFLX, AMZN, and MSFT for the originally chosen period (6 October 2022 to 10 April 2023) by Yu et al. (2023). In addition to reproducing their results, we incorporate a broader set of benchmarks, including traditional rule-based, ML, and DL strategies. Since prior studies often omit critical details such as the exact risk-free rate and transaction costs, both of which are essential for realistic backtesting, we explicitly set a risk-free rate of 0.03 based on historical average rates, and adopt Moomoo’s [^2] https://www.moomoo.com/ca/support/topic10_122?lang=en-us standard US commission fee (\$0.0049/share, minimum \$0.99/order), which is widely comparable to other brokerage platforms, such as HSBC and TradeUp [^3] https://www.tradeup.com/pricing/detail.
Symbol Type Strategy SPR $\uparrow$ CR $\uparrow$ MDD $\uparrow$ AV $\downarrow$
FinMem Selection (2022-10-06 to 2023-04-10)
TSLA Rule Based Buy and Hold -0.342 -20.483 -52.729 55.910
SMA Cross -0.293 -5.540 -18.517 38.602
WMA Cross 0.215 3.741 -18.492 42.062
ATR Band -0.595 -19.142 -39.599 42.161
Bollinger Bands -0.769 -24.747 -44.655 45.366
Turn of The Month 0.219 3.639 -11.642 31.042
Predictor ARIMA 0.601 15.007 -24.446 41.402
XGBoost 0.331 6.213 -35.374 37.729
RL A2C - - - -
DDPG - - - -
PPO -0.343 -20.425 -52.594 55.861
SAC - - - -
TD3 - - - -
LLM FinMem (GPT-4o-mini) 0.593 10.520 -30.144 47.518
FinMem (GPT-4o) 0.587 10.710 -33.856 53.156
FinMem (reported) 2.679 61.776 -10.800 46.865
FinAgent - - - -
NFLX Rule Based Buy and Hold 1.326 43.079 -20.184 41.523
SMA Cross -1.020 -8.285 -15.942 20.477
WMA Cross -0.803 -6.004 -14.290 19.826
ATR Band 0.150 2.992 -12.231 19.314
Bollinger Bands -0.558 -4.996 -13.244 16.754
Turn of The Month 0.559 8.383 -10.641 17.194
Predictor ARIMA 1.159 23.783 -15.043 25.749
XGBoost 0.770 10.134 -11.246 14.928
RL A2C - - - -
DDPG - - - -
PPO - - - -
SAC - - - -
TD3 1.325 42.872 -20.121 41.448
LLM FinMem (GPT-4o-mini) -2.278 -2.460 -2.460 3.476
FinMem (GPT-4o) 1.543 41.167 -20.417 51.030
FinMem (reported) 2.017 36.449 -15.850 36.434
FinAgent 1.543 41.167 -20.417 51.030
AMZN Rule Based Buy and Hold -0.460 -13.250 -31.546 35.624
SMA Cross -0.420 -4.433 -18.910 27.084
WMA Cross -0.563 -6.121 -21.030 26.831
ATR Band 0.622 11.007 -15.842 23.272
Bollinger Bands -0.402 -7.105 -20.615 26.559
Turn of The Month -0.037 0.039 -14.892 20.722
Predictor ARIMA -0.225 -4.752 -20.046 26.899
XGBoost 1.955 42.468 -8.816 25.135
RL A2C -0.462 -13.186 -31.397 35.496
DDPG -0.462 -13.186 -31.397 35.496
PPO -0.576 -9.485 -22.761 24.169
SAC - - - -
TD3 -0.462 -13.186 -31.397 35.496
LLM FinMem (GPT-4o-mini) 0.589 4.257 -2.746 9.875
FinMem (GPT-4o) -0.610 -13.764 -30.681 40.316
FinMem (reported) 0.233 4.885 -22.929 42.658
FinAgent -2.444 -3.819 -3.819 4.374
MSFT Rule Based Buy and Hold 0.974 21.171 -14.192 28.327
SMA Cross 1.515 18.289 -8.746 20.821
WMA Cross 1.334 16.576 -8.883 21.503
ATR Band 1.036 12.979 -7.709 15.005
Bollinger Bands 2.115 31.619 -3.475 18.243
Turn of The Month -0.034 0.970 -11.955 15.097
Predictor ARIMA 2.245 44.777 -7.121 22.636
XGBoost 0.895 12.678 -10.734 16.721
RL A2C 0.973 21.059 -14.120 28.117
DDPG - - - -
PPO - - - -
SAC - - - -
TD3 - - - -
LLM FinMem (GPT-4o-mini) -0.086 -1.679 -14.963 28.246
FinMem (GPT-4o) 0.307 -3.953 -45.462 89.114
FinMem (reported) 1.440 23.261 -14.989 32.562
FinAgent 0.738 11.972 -17.499 34.672
COIN Rule Based Buy and Hold 0.170 -12.729 -54.402 86.718
SMA Cross 0.566 15.215 -33.219 90.853
WMA Cross 0.517 12.655 -33.259 94.180
ATR Band 0.933 25.169 -22.906 38.716
Bollinger Bands -0.795 -24.371 -40.733 43.904
Turn of The Month 0.713 21.242 -21.522 49.357
LLM FinMem (GPT-4o-mini) -1.674 -8.196 -8.196 11.582
FinMem (GPT-4o) -0.741 -39.669 -51.436 88.508
FinMem (reported) 0.717 34.983 -35.753 89.752
FinAgent 0.306 -8.605 -55.827 104.742
Table 2: Backtest performance of benchmark strategies over the selective period (2022-10-06 to 2023-04-10), as presented in Yu et al. (2023). “-” metrics indicate no trading activities were triggered. Top and second-best values are marked in red and blue.
Table 2 summarises these results. Our analysis reveals that LLM investors are not universally superior, even within their originally favoured setups. Specifically, FinMem shows significant outperformance only for TSLA, whereas traditional benchmarks remain competitive or superior for other symbols. These observations caution against overly optimistic interpretations from selective evaluations. Additionally, we find that LLM-based strategies exhibit high annual volatility and substantial maximum drawdowns, indicating a high-risk profile. This observation underscores the importance of explicit risk assessments in evaluating such strategies.
Additionally, our results highlight that the choice of underlying language model dramatically impacts strategy performance. The original FinMem uses GPT-4; however, switching to GPT-4o or GPT-4o-mini significantly alters results. Interestingly, larger models (GPT-4o) do not consistently outperform smaller ones (GPT-4o-mini), suggesting that model size alone is not predictive of superior performance and that model selection itself may introduce data-snooping biases. Further evidence in Appendix D reinforces this instability, where even a slight two-month extension of the evaluation period results in substantial variation for LLM-based strategies.
Extending the Evaluation Period.
To further illustrate the limitations of short evaluation horizons, we extend the evaluation period (2004–2024), while keeping the same four symbols (TSLA, NFLX, AMZN, MSFT). This longer historical perspective aims to assess the true robustness of the previously reported LLM outperformance.
Symbol Type Strategy SPR $\uparrow$ STR $\uparrow$ AR $\uparrow$ MDD $\uparrow$ AV $\downarrow$
Selected 4 (2004-01-01 to 2024-01-01)
TSLA Rule Based Buy and Hold 0.631 0.915 37.767 -50.839 45.242
SMA Cross 0.011 0.106 16.352 -42.975 29.144
WMA Cross 0.218 0.375 19.743 -31.866 27.471
ATR Band 0.021 0.062 -0.277 -38.516 27.262
Bollinger Bands 0.193 0.294 4.279 -37.166 26.267
Trend Following 0.139 0.624 17.342 -13.508 14.750
Turn of The Month 0.207 0.352 7.868 -27.904 23.595
Predictor ARIMA 0.681 1.003 24.138 -30.450 27.612
XGBoost 0.142 0.370 10.877 -22.901 19.537
RL A2C 0.172 0.249 3.875 -27.367 22.890
DDPG 0.446 0.653 26.265 -21.626 21.892
PPO 0.469 0.663 28.189 -46.810 40.156
SAC 0.119 0.190 6.654 -11.042 9.902
TD3 0.417 0.604 23.336 -33.725 30.233
LLM FinMem 0.641 1.069 42.153 -34.234 35.030
FinAgent 0.206 0.649 38.591 -36.930 38.302
NFLX Rule Based Buy and Hold 0.622 0.952 23.919 -48.119 41.704
SMA Cross -0.126 0.015 1.219 -32.856 22.373
WMA Cross -0.091 -0.008 1.586 -34.378 23.275
ATR Band 0.118 0.273 1.018 -37.646 24.373
Bollinger Bands 0.075 0.381 0.284 -34.002 23.090
Trend Following -0.248 -0.094 3.433 -15.066 14.528
Turn of The Month 0.288 0.489 7.122 -21.646 17.168
Predictor ARIMA 0.659 1.035 19.022 -27.567 25.514
XGBoost 0.202 0.355 4.957 -21.301 17.302
RL A2C 0.171 0.243 4.359 -20.960 16.129
DDPG 0.211 0.340 7.370 -27.922 22.317
PPO 0.541 0.814 19.279 -39.615 33.630
SAC 0.186 0.285 8.397 -9.545 9.216
TD3 0.291 0.431 10.900 -21.451 19.304
LLM FinMem 0.293 0.622 12.566 -27.721 26.876
FinAgent -0.419 0.621 22.543 -20.466 22.838
AMZN Rule Based Buy and Hold 0.551 0.829 15.997 -36.842 30.859
SMA Cross -0.265 -0.247 0.582 -24.203 17.142
WMA Cross -0.144 -0.100 2.281 -22.867 16.329
ATR Band 0.453 1.014 5.639 -19.922 15.166
Bollinger Bands 0.020 0.126 0.900 -23.757 15.764
Trend Following -0.041 0.438 7.153 -10.265 12.342
Turn of The Month -0.033 -0.013 1.479 -20.472 15.726
Predictor ARIMA 0.339 0.504 7.523 -20.612 19.115
XGBoost -0.587 -0.366 1.200 -13.659 11.106
RL A2C 0.165 0.247 3.925 -14.841 11.654
DDPG 0.252 0.391 7.059 -10.153 9.306
PPO 0.505 0.767 13.831 -29.128 24.392
SAC 0.179 0.257 4.438 -14.093 11.665
TD3 0.382 0.597 11.738 -21.942 19.149
LLM FinMem 0.188 0.340 5.695 -28.296 24.786
FinAgent 0.364 0.663 12.699 -25.516 25.390
MSFT Rule Based Buy and Hold 0.461 0.620 11.237 -25.457 21.787
SMA Cross -0.172 -0.167 1.948 -16.610 12.051
WMA Cross -0.226 -0.233 0.827 -16.708 12.083
ATR Band 0.316 0.636 5.719 -11.893 10.881
Bollinger Bands -0.053 -0.028 1.578 -16.102 11.933
Trend Following -0.366 -0.205 1.966 -8.846 7.714
Turn of The Month -0.264 -0.343 -0.180 -14.309 10.438
Predictor ARIMA 0.304 0.466 8.207 -15.227 13.819
XGBoost 0.171 0.322 5.890 -10.523 10.335
RL A2C 0.279 0.380 7.478 -13.447 11.933
DDPG 0.300 0.403 7.838 -14.007 12.351
PPO 0.344 0.463 8.589 -16.697 14.410
SAC 0.216 0.288 5.329 -14.866 11.835
TD3 0.050 0.070 1.405 -9.491 6.648
LLM FinMem 0.203 0.293 4.567 -19.270 17.891
FinAgent 0.285 0.432 11.123 -18.596 18.863
Table 3: Backtest performance of benchmark strategies for symbols from Yu et al. (2023) over extended period (2004-01-01 or earliest available to 2024-01-01). Top and second-best values are marked in red and blue.
Table 3 summarises these extended period results. Crucially, we observe that extending the evaluation horizon significantly diminishes the perceived superiority of LLM investors. Over two decades, traditional strategies like Buy and Hold consistently rank among the top performers across most symbols. For example, TSLA remains the only symbol where LLM investors (FinMem, FinAgent) retain clear superiority in annual return, whereas for NFLX, AMZN, and MSFT, Buy and Hold or other strategies outperform or match LLM strategies. This observation demonstrates that previously reported LLM advantages are likely short-lived, potentially hand-picked, and heavily dependent on the chosen evaluation period. Such fragility aligns with a central tenet in financial economics: asset prices are generally unpredictable in the long run, an assumption formalised in the Efficient Market Hypothesis (Fama, 1970). This foundational view cautions against over-interpreting transient outperformance as evidence of persistent edge, particularly when evaluation setups are narrow or selective.
It is crucial to note that we cannot yet conclude that benchmark strategies cannot outperform the market. As mentioned, backtesting only on popular stocks may inadvertently introduce survivorship bias, as these stocks have gained popularity due to past success during prolonged bull markets. Thus, expanding the range of symbols is essential to ensure a more systematic and unbiased evaluation.
6.2 Fair Comparisons with the Composite Approach
Selection Strategy Type Timing Strategy SPR $\uparrow$ STR $\uparrow$ AR $\uparrow$ MDD $\uparrow$ AV $\downarrow$
Random 5 S&P 500 (91 symbols) Rule Based Buy and Hold 0.315 0.456 6.694 -35.130 27.410
SMA Cross -0.298 -0.290 0.446 -22.292 15.774
WMA Cross -0.299 -0.305 0.232 -22.754 15.528
ATR Band 0.232 0.425 5.119 -21.535 16.113
Bollinger Bands 0.129 0.288 3.521 -22.487 16.290
Trend Following -0.389 -0.198 2.525 -8.587 8.223
Turn of The Month 0.015 0.072 2.870 -18.582 13.542
Predictor ARIMA 0.255 0.434 6.928 -21.691 17.504
XGBoost -0.055 0.028 3.089 -17.160 13.075
RL A2C 0.086 0.122 1.902 -9.220 6.887
DDPG 0.089 0.125 0.844 -15.185 11.624
PPO 0.179 0.256 3.282 -18.395 13.783
SAC 0.097 0.142 1.389 -16.058 12.375
TD3 0.173 0.248 3.682 -14.471 11.565
LLM FinMem -0.253 0.114 -0.094 -24.243 21.214
FinAgent 0.094 0.323 4.477 -28.059 26.387
Momentum Factor (84 symbols) Rule Based Buy and Hold 0.384 0.694 9.916 -32.596 37.421
SMA Cross -0.251 0.008 2.109 -19.438 20.050
WMA Cross -0.169 0.051 3.674 -18.651 20.330
ATR Band 0.197 0.595 4.314 -19.407 20.038
Bollinger Bands 0.114 0.702 1.881 -19.451 21.555
Trend Following 0.119 0.531 6.380 -15.726 18.696
Turn of The Month 0.056 0.662 3.197 -18.108 18.055
Predictor ARIMA 0.542 1.043 13.257 -18.277 22.892
XGBoost 0.094 1.525 6.131 -12.754 17.238
RL A2C 0.105 0.171 2.488 -14.452 14.815
DDPG 0.209 0.366 5.973 -14.255 16.499
PPO 0.185 0.308 1.939 -23.177 25.527
SAC 0.195 0.321 5.591 -12.235 16.144
TD3 0.186 0.293 3.464 -14.593 14.953
LLM FinMem 0.025 0.170 3.649 -23.335 28.078
FinAgent 0.104 0.534 13.950 -20.675 30.635
Volatility Effect (63 symbols) Rule Based Buy and Hold 0.703 1.291 7.898 -14.146 14.720
SMA Cross -0.568 -0.544 0.781 -9.296 8.665
WMA Cross -0.665 -0.348 1.908 -8.481 8.573
ATR Band -0.026 0.120 2.798 -8.032 7.951
Bollinger Bands -0.077 0.029 2.503 -7.618 7.774
Trend Following 0.230 0.619 5.503 -8.115 9.297
Turn of The Month -0.156 -0.095 2.881 -6.889 7.233
Predictor ARIMA 0.325 0.838 4.898 -9.111 9.807
XGBoost -0.108 -0.055 2.775 -6.676 7.077
RL A2C 0.421 0.795 4.620 -4.428 5.149
DDPG 0.373 0.725 3.768 -6.681 6.693
PPO 0.514 0.972 5.805 -8.757 9.461
SAC 0.402 0.810 3.527 -4.821 5.030
TD3 0.269 0.394 4.610 -5.442 5.992
LLM FinMem -0.228 0.483 4.061 -10.860 11.641
FinAgent 0.241 0.527 4.954 -10.268 11.502
Table 4: Backtest performance under the Composite setup, using three different selection strategies across historical S&P 500 constituents (2004–2024), including delisted symbols. Top and second-best values are marked in red and blue.
To overcome the aforementioned biases, we introduce the Composite evaluation setup within our FINSABER framework. This setup integrates systematic selection-based strategies to expand and diversify the stock universe, explicitly addressing survivorship and data-snooping biases. Specifically, we use three unbiased stock selection approaches (details in Appendix B.2): Random Five, Momentum Factor (Muller and and, 2010), and Volatility Effect (Blitz and Vliet, 2007).
To effectively mitigate survivorship bias, our selection process uses only historical constituent lists of the S&P 500 available at the beginning of each evaluation period and explicitly includes delisted symbols. This ensures we avoid inadvertently selecting stocks that have survived due to their sustained performance over time. Moreover, the data-snooping bias is mitigated by evaluating a large and diversified universe of symbols (91, 84, and 63 symbols for Random Five, Momentum-based, and Volatility-based selection, respectively) over an extensive period of two decades, thus eliminating the biases associated with cherry-picking favourable stocks or short evaluation windows. These numbers refer to the total distinct symbols encountered across all rolling windows during evaluation, where we reselect stocks in each window.
Table 4 summarises these comprehensive evaluations. Results obtained through this unbiased and systematic approach further validate our previous findings from the selected-four evaluation. Specifically, both the Random Five and Momentum-based selections reinforce the conclusion that the previously claimed superiority of LLM investors is largely driven by selective evaluation setups. For instance, in the Random Five setup, Buy and Hold, ATR Band and ARIMA outperform FinMem and FinAgent in terms of risk-adjusted metrics. Similarly, ARIMA and simple rule-based strategies often perform better than LLM-based methods under the Momentum-based selection. In the Volatility-based selection, traditional methods dominate even more clearly: Buy and Hold achieves the highest Sharpe (0.703), stability (1.291), and annual return (7.898%), while PPO and ARIMA again show strong all-round performance. LLM-based methods lag behind, with FinAgent offering moderate returns but lower Sharpe (0.241) and larger drawdowns. Notably, our reported LLM performances do not adjust for potential data leakage: given the use of pretrained models like GPT-4o, the LLMs may have seen parts of the data during training—yet they still fail to outperform traditional strategies under fair evaluation, casting further doubt on their real-world advantage.
Nevertheless, it is important to acknowledge that LLM-based strategies still show potential regarding absolute annual returns. For instance, FinAgent achieves the highest annual return (13.950%) in the Momentum-based selection setup. However, the relatively weaker performance observed in Sharpe ratio (0.104) and maximum drawdown metrics suggests a clear need for improved risk management within LLM-driven approaches before they can be reliably adopted in practice.
Moreover, by comparing Buy and Hold with different selection strategies, we clearly identify the relative effectiveness of each selection strategy: Volatility selection (Sharpe 0.703) outperforms Momentum (0.384), which in turn surpasses Random Five (0.315). RL-based methods exhibit the clearest alignment with selection quality. Strategies like PPO, SAC, and TD3 systematically achieve their best performance under the Volatility selection and degrade under the other two. This suggests RL methods are more sensitive to the quality of the stock candidates. Among LLM strategies, FinAgent exhibits a greater dependency on selection quality than FinMem.
Overall, these results not only confirm our earlier insights but also underscore the critical importance of unbiased, systematic stock-selection methodologies for accurately assessing the true capabilities and limitations of LLM-based investing strategies.
7 Market Regime Analysis
Figure 2: Average Sharpe ratio by regime for all benchmarking strategies. Green = strong, red = weak.
A remaining question in evaluating LLM-based investing strategies is understanding whether strategies adapt appropriately across varying market conditions. As Kim et al. (2011) demonstrated, financial markets exhibit time-varying predictability and uncertainty across different economic, financial, and political regimes. Some strategies may be able to exploit these variations, while others may struggle to adapt effectively. Different market environments including bull, bear, and sideways, pose distinct challenges and opportunities: bull markets reward aggressive positioning and substantial exposure, bear markets require risk management, and sideways markets test strategies’ ability to navigate uncertainty without clear trends. By decomposing performance into these regimes, we can identify whether strategies are overly conservative, missing opportunities during bullish periods, or excessively aggressive, suffering large losses during downturns. Clarifying these regime-specific behaviours is crucial for interpreting the strengths and weaknesses of LLM-based investing strategies and for informing their future development (Hui and Chan, 2018).
We label each calendar year based on the annual return of the S&P 500: $R_{y}=\frac{P_{T}-P_{0}}{P_{0}}$, where $P_{0}$ and $P_{T}$ are the adjusted closing prices on the first and last trading days of year $y$. A year is classified as bull if $R_{y}\geq+20\%$, bear if $R_{y}\leq-20\%$, and sideways otherwise. The ±$\pm 20\%$ threshold follows standard industry convention (Zweig, 2019).
To analyse regime-specific performance, we use our composite setup under three selection strategies: Random 5, Momentum Factor (Muller and and, 2010), and Volatility Effect (Blitz and Vliet, 2007). For each timing strategy, we retrieve the Sharpe ratio within each 1-year window from Table 4. These are then averaged per {strategy, regime} pair to produce stable performance indicators across market conditions. Figure 2 illustrates the results, with green indicating strong Sharpe ratio and red signifying the opposite.
Traditional rule-based and predictor-based methods still set the standard. ATR Band, Turn of the Month and ARIMA deliver positive Sharpe in every regime, while Buy and Hold, our passive yardstick, posts 0.61 in bulls, 0.48 in sideways markets and only -0.28 in bears. No active strategy surpasses this passive Sharpe ratio in the bull regime, suggesting that many strategies, including the LLM ones, may struggle to fully capitalise on strong up-trends.
RL algorithms sit in the middle. A2C and DDPG pick up part of the upside and limit losses; PPO and SAC swing with volatility and underperform ARIMA once conditions turn.
The LLM strategies perform poorly. FinAgent records Sharpe 0.12 in bulls and -0.38 in bears; FinMem slides from -0.19 to -0.97. Both are too cautious when risk is rewarded and too aggressive when it is penalised. FinAgent is the better of the two, halving the bear-market shortfall relative to Buy and Hold and keeping a small positive Sharpe in neutral conditions, but it still trails every rule-based or predictor benchmark.
These results suggest two directions for future LLM investors. A first direction concerns improving trend-detection so the strategy can at least match passive equity beta in up-markets. A second direction relates to embedding explicit regime-aware risk controls that scale exposure down as volatility or draw-down risk rises. Balancing risk-taking and risk management, aggression and defence, rather than increasing model size, appears the key to closing the gap with traditional methods.
8 Conclusion
We reassess the robustness of LLM timing-based investing strategies using FINSABER, a comprehensive framework that mitigates backtesting biases and extends both the evaluation horizon and symbol universe. Our results show that the perceived superiority of LLM-based methods deteriorates under more robust and broader long-term testing. Regime analysis further reveals that current strategies miss upside in bull markets and incur heavy losses in bear markets due to poor risk control.
We identify two priorities for future LLM-based investors: (1) enhancing uptrend detection to match or exceed passive exposure, and (2) including regime-aware risk controls to dynamically adjust aggression. Addressing these dimensions rather than increasing framework complexity is the key to building practical, reliable strategies.
Finally, our cost analysis (Appendix F) shows that large-scale LLM backtesting is financially intensive. Future work should pursue cost-efficient model designs and incorporate API costs into performance evaluation.
Limitations
There are several limitations to our current study. First, we did not individually tune the traditional rule-based strategies for each rolling evaluation window. Typically, applying domain-specific market insights to optimise parameters can significantly enhance the performance of these methods. However, we argue that our current configuration remains valid and effectively demonstrates the competitive disadvantage faced by LLM strategies. Indeed, tuning the parameters of traditional rule-based strategies would likely elevate their performance further, reinforcing rather than undermining our main conclusions.
Second, our evaluation has not fully eliminated look-ahead bias. Pre-trained LLMs, due to their inherent training corpus, may inadvertently contain stock-related information from historical periods overlapping our test sets. Despite this potential data leakage, the observed underperformance of LLM strategies strengthens our critical assessment. Explicitly addressing this look-ahead concern through controlled model training or careful exclusion of financial data from training corpora will be an important avenue for future research.
Third, to ensure experiment reproducibility, we restricted our analysis to publicly available data, excluding proprietary sources such as private newsfeeds, earning transcripts, or expert analyses. Nonetheless, the FINSABER framework was deliberately designed to be modular and extensible, allowing researchers with access to private data to easily integrate additional information sources. Our primary goal remains providing a rigorous, long-term evaluation pipeline that minimises selective reporting. Researchers lacking proprietary data can fully replicate our results using openly accessible resources.
Appendix A Data Collection
Our multi-source data comprises daily stock prices, daily financial news, and 10-Q and 10-K filings.
Daily Stock Prices.
We collect daily price data for over 7,000 U.S. equities spanning from 2000 to 2024. In addition to actively traded stocks, our dataset includes delisted symbols that were historically part of the S&P 500 index, based on the archived constituent list [^4] https://github.com/fja05680/sp500. This inclusion enhances the historical completeness of our dataset and mitigates survivorship bias within the context of index-based evaluations.
Financial News.
The financial news dataset, initially compiled by Dong et al. (2024), comprises 15.7 million records pertaining to 4,775 S&P 500 companies, spanning the years 1999 to 2023. We have organised the news by aligning it with the respective companies and indexing it by date.
10K & 10Q Filings.
We collect 10-K and 10-Q filings for companies included in the Russell 3000 index, sourced from the US Securities and Exchange Commission (SEC) EDGAR database. These filings are publicly available and accessed via the SEC-API [^5] https://sec-api.io/, which allows programmatic retrieval and parsing. We preprocess the HTML documents and segment them into standardized sections, such as Risk Factors, MD&A, and Financial Statements, to support fine-grained analysis. Each filing is indexed by company identifier and filing date to enable alignment with other datasets.
Extensibility.
All datasets used in this framework can be seamlessly substituted with proprietary or higher-resolution alternatives if available. Researchers may incorporate paid datasets such as premium financial news (e.g., Alpaca Markets [^6] https://alpaca.markets/, Refinitiv [^7] https://www.lseg.com/en), earnings call transcripts, analyst research reports, or other modalities including video or audio. Integration is supported through the implementation of a custom dataset class, allowing modular and flexible replacement of any data stream within the pipeline.
Appendix B FinSABER Strategies Base
B.1 Timing-based Strategies
Open-Source LLM investors.
This category includes FinMem (Yu et al., 2023) and FinRobot (Yang et al., 2024). We acknowledge other works, such as FinAgent (Zhang et al., 2024b), FinCon (Yu et al., 2024) but they are not (yet) open-source, which prevents us from generating backtesting results.
Traditional Rule-Based (Indicator-Based) Strategies.
We implement and cover several well-known traditional rule-based (indicator-based) investing strategies, such as Buy and Hold, Simple Moving Average Crossover, Weighted Moving Average Crossover, ATR Band, Bollinger Bands (Bollinger, 2002), Trend Following (Wilcox et al., 2005), and Turn of the Month (McConnell and Xu, 2008). These strategies typically rely on one or multiple technical indicators or domain-based rules to generate timely buy/sell signals, aiming to exploit identifiable market patterns or anomalies.
It is noteworthy that traditional strategies are often overlooked, with many existing works focusing solely on Buy and Hold. However, other established strategies listed above have also endured over time and demonstrated their effectiveness.
ML/DL Forecaster-Based Strategies
In contrast to fixed rules or indicator-based triggers, these strategies rely on data-driven models—often statistical or neural network forecasters—to predict future price movements. Specifically, they buy or hold if an uptrend is indicated and sell (or go short) otherwise. This can be viewed as a relatively naive application of ML/DL forecasters, but it is widely used as a benchmark method for such models. Although one could consider the forecast output as a type of “indicator”, the reliance on predictive algorithms capable of uncovering complex patterns sets these methods apart from purely rule-based approaches. We include the well-known ARIMA (Box and Jenkins, 1990) and XGBoost (Chen and Guestrin, 2016) in this category and also cover forecasters based on LLMs, differentiating them from the LLM investors.
RL-Based Strategies
Additionally, we incorporate RL-based strategies to evaluate their performance in investing environments. Specifically, we implement Advantage Actor-Critic (A2C), Proximal Policy Optimization (PPO), Deep Deterministic Policy Gradient (DDPG), Twin Delayed Deep Deterministic Policy Gradient (TD3), and Soft Actor-Critic (SAC), leveraging the FinRL framework (Liu et al., 2022, 2024). These RL algorithms are widely used in financial markets due to their ability to learn optimal investing policies from historical data and adapt to evolving market conditions. A2C and PPO, as policy-gradient methods, optimise investing decisions while ensuring stability, whereas DDPG, TD3, and SAC extend this capability to continuous action spaces, rendering them well-suited for asset allocation and portfolio optimisation tasks.
B.2 Selection-based Strategies
(1) Random Five, which involves randomly selecting five stocks from historical S&P 500 constituents available at each rolling evaluation period’s start date, (2) Momentum Factor Selection, which selects stocks based solely on past returns at each period’s start date, following a momentum factor-based strategy (Muller and and, 2010), and (3) Volatility Effect Selection, which selects stocks with the lowest historical return volatility, capturing the empirically observed inverse relationship between volatility and risk-adjusted performance (Blitz and Vliet, 2007).
Appendix C Evaluation Metrics
We group evaluation metrics into three categories, each targeting a distinct aspect of strategy performance:
C.1 Return Metrics
Annualised Return (AR).
Measures the geometric average yearly return: $R_{\textrm{annual}}=(1+C)^{\frac{252}{T}}-1$ where $C=\prod_{t=1}^{T}(1+R_{t})-1$.
Cumulative Return (CR).
Total portfolio return over the test period: $C=\prod_{t=1}^{T}(1+S_{t}\cdot R_{t})-1$, where $S_{t}$ is the position and $R_{t}$ is the market return.
C.2 Risk Metrics
Annualised Volatility (AV).
Volatility of daily returns, scaled annually: $\sigma_{\textrm{annual}}=\sigma_{\textrm{daily}}\times\sqrt{252}$.
Maximum Drawdown (MDD).
Largest drop from peak portfolio value: $\textrm{MDD}=-\max_{t\in[1,T]}\left(\frac{\textrm{Peak}_{t}-V_{t}}{\textrm{Peak}_{t}}\right)$.
C.3 Risk-adjusted Performance Metrics
Sharpe Ratio (SPR).
Return per unit of total volatility: $\textrm{SPR}=\frac{\overline{R_{t}}-\frac{r_{f}}{252}}{\sigma_{\textrm{daily}}}\times\sqrt{252}$.
Sortino Ratio (STR).
Return per unit of downside deviation: $\textrm{STR}=\frac{\overline{R_{t}}-\frac{r_{f}}{252}}{\sigma_{\textrm{down}}}\times\sqrt{252}$.
Where $\overline{R_{t}}$ is the average daily return, $r_{f}$ is the annual risk-free rate, and $\sigma_{\textrm{down}}$ is the standard deviation of negative returns.
Appendix D Extra Results on Selective Symbols
Tables 2 and 5 further substantiate our findings by highlighting the performance instability of FinMem and FinAgent when extending evaluation periods even marginally. Specifically, extending the evaluation by just two months beyond the originally reported periods (Yu et al., 2023) results in notable inconsistencies in critical performance metrics. It should be noted that the results for the LLM strategies are retrieved from Yu et al. (2024), while the traditional rule-based results presented are based on our implementations.
For instance, FinMem exhibited a drastic change in cumulative returns for MSFT from a reported 23.261% down to -22.036%, and a reduction in Sharpe ratios from 1.440 to -1.247. Similarly, for NFLX, the Sharpe ratio for FinMem shifted dramatically from a reported 2.017 to -0.478. These examples underscore the sensitivity of LLM-based investing strategies to minor shifts in market conditions and reinforce our argument about the necessity of comprehensive and temporally robust evaluations to accurately assess the reliability and generalisability of these models.
TSLA AMZN NIO MSFT
Type Strategy SPR CR MDD SPR CR MDD SPR CR MDD SPR CR MDD
FinCon Selection (2022-10-05 to 2023-06-10)
Buy And Hold 0.247 2.056 -54.508 0.150 2.193 -32.177 -0.858 -51.569 -53.563 1.071 32.629 -14.452
SMA Cross -0.151 -3.973 -23.173 0.599 13.731 -18.910 0.810 22.047 -17.976 1.641 32.057 -8.746
WMA Cross 1.104 32.058 -18.492 0.513 11.765 -21.030 -0.771 -9.412 -18.732 1.526 30.344 -8.883
ATR Band -0.554 -22.136 -39.599 0.494 11.007 -15.842 0.681 24.684 -21.229 0.827 12.979 -7.709
Bollinger Bands -0.249 -12.756 -44.655 -0.381 -7.105 -20.615 0.940 25.476 -16.623 1.759 31.619 -3.475
Rule Based Turn of The Month 0.928 27.850 -11.642 0.123 3.487 -14.892 0.874 31.344 -17.995 0.407 7.744 -11.955
FinGPT 0.044 1.549 -42.400 -1.810 -29.811 -29.671 -0.121 -4.959 -37.344 1.315 21.535 -16.503
FinMem 1.552 34.624 -15.674 -0.773 -18.011 -36.825 -1.180 -48.437 -64.144 -1.247 -22.036 -29.435
FinAgent 0.271 11.960 -55.734 -1.493 -24.588 -33.074 0.051 0.933 -19.181 -1.247 -27.534 -39.544
LLM FinCon 1.972 82.871 -29.727 0.904 24.848 -25.889 0.335 17.461 -40.647 1.538 31.625 -15.010
AAPL GOOG NFLX COIN
Type Strategy SPR CR MDD SPR CR MDD SPR CR MDD SPR CR MDD
FinCon Selection (2022-10-05 to 2023-06-10)
Buy And Hold 0.906 24.558 -19.508 0.683 20.884 -20.278 1.594 77.367 -20.421 0.024 -23.761 -54.402
SMA Cross 1.423 21.054 -6.030 0.382 8.497 -17.035 -0.855 -8.393 -18.545 0.232 1.286 -35.559
WMA Cross 1.648 25.257 -6.114 0.635 13.659 -14.985 -1.009 -9.479 -18.531 0.087 -7.461 -40.883
ATR Band 0.241 4.522 -5.159 0.067 2.616 -13.522 0.522 10.739 -12.231 0.777 25.169 -22.906
Bollinger Bands - - - 0.365 7.526 -13.522 -0.182 -0.710 -13.244 -0.705 -24.371 -40.733
Rule Based Turn of The Month 0.098 3.337 -12.498 0.343 7.188 -13.519 0.987 18.942 -10.641 -0.020 -8.999 -33.895
FinGPT 1.161 20.321 -16.759 0.011 0.242 -26.984 0.472 11.925 -20.201 -1.807 -99.553 -74.967
FinMem 0.994 12.397 -11.268 0.018 0.311 -21.503 -0.478 -10.306 -27.692 0.017 0.811 -50.390
FinAgent 1.041 20.757 -19.896 -1.024 -7.440 -10.360 1.960 61.303 -20.926 -0.106 -5.971 -56.882
LLM FinCon 1.597 27.352 -15.266 1.052 25.077 -17.530 2.370 69.239 -20.792 0.825 57.045 -42.679
Table 5: Backtest performance of traditional rule-based (indicator-based) strategies and FinCon over the selective period (2022-10-05 to 2023-06-10), as presented in Yu et al. (2024), evaluated using four metrics: cumulative return (CR), Sharpe ratio (SPR), annual volatility (AV), and maximum drawdown (MDD). The best metrics are highlighted in red, while the second best are marked in blue. “-” metrics across the board indicate no trade signals were triggered.
Appendix E Technical Details
FINSABER Implementation.
The backtesting framework and traditional rule-based strategies in FINSABER are implemented using BackTrader [^8] https://www.backtrader.com/ and Papers With Backtest [^9] https://paperswithbacktest.com/. Reinforcement learning-based methods are implemented using FinRL (Liu et al., 2021). FINSABER supports two operational modes: “LLM” mode and “BT” mode. The “LLM” mode is tailored for strategies that leverage multi-modal inputs, including financial news and regulatory filings. In contrast, the “BT” mode is built directly on BackTrader, offering robust support for traditional rule-based strategies while maintaining a familiar interface to facilitate easy migration from standard BackTrader workflows.
Experiment Rolling Windows.
We apply a rolling-window evaluation setup to ensure temporal robustness and reduce data-snooping bias. For the Selected 4 evaluation, we use a 2-year rolling window with a 1-year step, and allow strategies to use up to 3 years of prior data for training. For the Composite setup, we adopt a more frequent rebalancing scheme with a 1-year rolling window and a 1-year step, allowing up to 2 years of prior data. This adjustment reflects the observation that rebalancing every two years may be too infrequent to capture changing market dynamics. All experiments span the benchmark period from 2004 to 2024.
Parameters of Strategies.
Table 6 summarises the key hyperparameters used for each benchmark strategy in our experiments. These settings are largely drawn from standard defaults commonly used in the public implementations. For traditional rule-based strategies, optimal parameter selection often requires domain expertise or practitioner experience. Our goal is not to optimise each strategy’s absolute performance, but to provide a fair and consistent baseline under a unified evaluation framework. We encourage future researchers to explore parameter optimisation techniques (e.g., grid search, Bayesian tuning) if desired.
Strategies Parameters
SMA Cross short_window=10, long_window=20
WMA Cross short_window=10, long_window=20
ATR Band atr_period=14, multiplier=1.5
Bollinger Band period=20, devfactor=2.0
Trend Following atr_period=10, period=20
Turn of the Month before_end_of_month_days=5, after_start_of_month_business_days=3
ARIMA order=(5,1,0)
XGBoost num_boost_round=10, n_estimators=1000
RL total_timesteps=50000
FinMem model=gpt-4o-mini, top_k=3, embedding_model=text-embedding-ada-002, chunk_size=5000
FinAgent model=gpt-4o-mini, trader_preference=aggressive_trader, top_k=5, previous_action_look_back_days=14
Table 6: Default parameter settings for benchmark strategies.
Appendix F LLM Strategies Cost Analysis
To better understand the practical deployment of LLM-based investing strategies, we monitor the API costs associated with running backtests on the Composite experiment with Volatility Effect selection as a representative example. The cost for backtesting FinAgent was \$198.24, while FinMem incurred a significantly lower cost of \$31.79 using GPT-4o mini. This reflects the higher prompt complexity and more frequent calls involved in FinAgent’s multi-agent decision-making process.
Extrapolating from these numbers, we estimate that completing all Composite experiments required approximately \$700 in LLM API costs. The Selected 4 setup likely incurred even greater cost, given its larger rolling window size and the increased volume of financial news associated with these selectively popular symbols.
FinAgent was roughly 6 times more expensive than FinMem in our tests. Importantly, these figures only account for LLM generation costs (i.e., chat/completions endpoints), and do not include the cost of generating embeddings (e.g., via text-embedding-ada-002), which would further increase the total budget.
This observation raises a practical consideration for future research: when evaluating LLM-driven strategies, computational cost should be factored into the financial metrics, particularly for real-world deployment scenarios. Incorporating API usage cost into risk-adjusted performance metrics (e.g., Sharpe or Sortino) could provide a more holistic picture of strategy efficiency.
Recommendation.
For researchers with limited budget, we recommend adopting open-source LLMs (e.g., LLaMA, Qwen, Mistral) for benchmarking and prototyping. These models can be deployed locally or via cost-effective cloud infrastructure, significantly reducing evaluation costs while enabling reproducible experimentation.